OpenStudy (anonymous):

Find a model for simple harmonic motion if the position at t = 0 is 4, the amplitude is 4 centimeters, and the period is 6 seconds. A. d=4cos((π/3)t) B. d=4sin((π/3)t) C. d=4sin(6t) D. d=6cos(4t)

OpenStudy (anonymous):

2pi/a=6 a/2pi=1/6 a=2pi/6 a=pi/3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks!